# Parity superselection obstructs monogamy of mutual information in free fermions

### 14 Mar 2026 v1

@@ Abstract
We prove that free fermions in the spin (tensor product) factorization violate monogamy of mutual information: $I_{3}^{\mathrm{spin}}>0$ for three adjacent strips of width $w=1,2$ at all Fermi momenta, and for all $w$ at $z=k_{F}w<z^{*}\approx 1.329$. The proof rests on an exact operator identity—the fermionic and spin reduced density matrices of disjoint regions differ by the parity insertion $(-1)^{N_{B}}$ in the partial trace—and a rigorous entropy bound.
DMRG calculations on the $t$-$V$ chain quantify the effect for interacting fermions: the factorization contribution to the apparent $K$-dependence of $I_{3}$ exceeds the genuine interaction contribution by a factor of 8 at moderate filling, and accounts for ${\sim}80\%$ of the deviation observed in spin-basis numerics. Strong repulsion ($K\lesssim 0.7$) restores monogamy in both algebras.
@@ 1 Introduction
quantifies irreducible tripartite correlations [1]. States satisfying $I_{3}\leq 0$ for all subsystems obey monogamy of mutual information (MMI)—a property guaranteed by holographic duality [1, 2] but violated by free field theories [3, 4].
For lattice fermions, entanglement depends on the choice of operator algebra [5, 6, 7, 8]. Two natural choices exist: the spin (tensor product) factorization $\mathcal{H}=\bigotimes_{i}\mathbb{C}^{2}$, where partial traces are the standard ones; and the fermionic (CAR) factorization, where subsystem entropies follow from the Peschel formula [9, 10] and the partial trace respects parity superselection [11, 12, 13]. For contiguous regions the two coincide. For disjoint regions they do not.
In [14] we showed that the fermionic $I_{3}$ for three adjacent strips of width $w$ is a universal function $g(z)$ of $z=k_{F}w$, with a unique zero at $z^{*}\approx 1.329$. In this paper we prove that switching to the spin factorization changes the sign of $I_{3}$: the quantity $I_{3}^{\mathrm{spin}}$ is strictly positive for $w=1,2$ at all $z$, and for all $w$ at $z<z^{*}$ (Fig. 1).
Figure 1: (a) The fermionic $I_{3}^{\mathrm{ferm}}=g(z)$ (blue) changes sign at $z^{*}\approx 1.329$, while $I_{3}^{\mathrm{spin}}$ (green) remains positive for all $z$. The difference $\Delta S_{AD}$ (orange dashed) grows monotonically and overwhelms $|g|$ for $z>z^{*}$. (b) For $z>z^{*}$: the Gaussian budget $\mathcal{B}(z)$ (Theorem 1, purple) exceeds $|g(z)|$ (red shaded) with minimum margin $2.9\times$. Data for $w=2$ (representative; $w=1,3$ show the same qualitative behavior with margins in Table 1).
@@ 2 Setup
This is a special case of the general identity $I_{m}=(-1)^{m+1}\Delta^{m}S$ for $m$-partite information, decomposed into contiguous and non-contiguous parts. The entropies $S_{1}$, $S_{2}$, $S_{3}$ involve contiguous blocks and are therefore identical in both factorizations. The entire algebra dependence resides in $S_{AD}$.
@@ 3 The superselection defect identity
Equation (6) is equivalently a block decomposition in the $D$-parity basis: the $D$-parity-preserving blocks of $\rho^{\mathrm{ferm}}$ equal those of $\rho^{\mathrm{spin}}$, while the $D$-parity-changing blocks equal those of $\rho^{\mathrm{tw}}$. The defect $\rho^{\mathrm{spin}}-\rho^{\mathrm{tw}}$ measures the difference between the two factorizations. When $\rho^{\mathrm{tw}}=\rho^{\mathrm{spin}}$ (no parity interference), the defect vanishes and $\rho^{\mathrm{ferm}}=\rho^{\mathrm{spin}}$—the two algebras agree. The identity holds for any state (Gaussian or interacting, pure or mixed, equilibrium or not) and for any widths of $A$, $B$, $D$; in particular, it applies at finite temperature, where both $I_{3}$ and $\Delta S_{AD}$ are suppressed by thermal fluctuations but their ratio remains $O(1)$ as long as $T\ll E_{F}$. Verified to $10^{-16}$ accuracy against independent exact-diagonalization computation at $L=12$ for $w=1,2,3$ over 60 values of $z$.
@@
Coherence damping. For $D$-parity-changing elements, the parity insertion produces destructive interference: $\rho_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}X_{b}$ while $\rho_{ij}^{\mathrm{tw}}=\sum_{b}(-1)^{N_{b}}p_{b}X_{b}$. Numerically, $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ for all tested parameters ($w=1,2,3$, full range of $z$; also verified for the interacting $t$-$V$ chain at all tested couplings).
Entropy ordering. Since $\rho^{\mathrm{ferm}}$ differs from $\rho^{\mathrm{spin}}$ only in the $D$-parity-changing sector, and these elements are damped ($|\rho^{\mathrm{tw}}|\leq|\rho^{\mathrm{spin}}|$ numerically), the fermionic RDM is “more diagonal” than the spin one in the $D$-parity basis. This is the physical origin of $\Delta S_{AD}\geq 0$.
@@ 4 Entropy bound
Step 2 (Maximum entropy [16]). Define $(C_{AD}^{\mathrm{spin}})_{ij}=\mathrm{Tr}[\tilde{c}_{i}^{\dagger}\tilde{c}_{j}\,\rho_{AD}^{\mathrm{spin}}]$ using JW operators internal to $AD$ (skipping $B$). The diagonal blocks $C_{AA}$, $C_{DD}$ equal those of $C^{\mathrm{ferm}}$ (both regions contiguous). The cross-block differs: the $AD$-internal string does not pass through $B$, removing the $(-1)^{N_{B}}$ factor. Since $\rho_{AD}^{\mathrm{spin}}$ is non-Gaussian [11]: $S_{AD}^{\mathrm{spin}}\leq S_{G}(C_{AD}^{\mathrm{spin}})$.
@@
The bound is conservative: the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$ makes $\Delta S_{AD}=\mathcal{B}+\varepsilon>\mathcal{B}$ (numerically $\varepsilon$ is 0.1–3% of $S^{\mathrm{spin}}$). As a check: $\Delta S_{AD}\leq\ln 2$; our data satisfy $\Delta S_{AD}/\ln 2<0.21$.
@@
Immediate consequence. For $z<z^{*}$: $g(z)>0$ [14] and $\Delta S_{AD}\geq 0$, so $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}>0$ for all $w$. This is fully rigorous.
For $z>z^{*}$: $g(z)<0$, and we need $\Delta S_{AD}>|g(z)|$. Since $\Delta S_{AD}\geq\mathcal{B}$, a sufficient condition is:
Conjecture 1 (Spectral inequality).
For all $z>z^{*}$ and $w\geq 1$: $\mathcal{B}(z)>|g(z)|$.
@@ 5.1 Analytical proof for $w=1$
The fermionic cross-correlator is $c_{2}$. The spin cross-correlator $s_{2}$ differs by the absence of $(-1)^{N_{B}}$ in the sum. From the Slater determinant overlap formula [9, 10]:
@@
All seven entropies entering $I_{3}^{\mathrm{spin}}$ are now explicit elementary functions. Evaluating on a grid of $10^{4}$ points with spacing $\Delta z=3.14\times 10^{-4}$ and second-derivative bound $|I_{3}^{\prime\prime}|<0.20$, the interpolation error is $(\Delta z)^{2}|I_{3}^{\prime\prime}|/8<2.5\times 10^{-9}$. The minimum is:
$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0557\quad\text{at}\quad z=\frac{\pi}{2},$ (11)
exceeding the error bound by a factor of $2\times 10^{7}$.
@@ 5.2 Certified computation for $w=2$
Grid of 500 points, spacing $\Delta z=6.1\times 10^{-3}$, curvature bound $|I_{3}^{\prime\prime}|<1.87$ (obtained by evaluating the numerical second derivative on a refined grid of $5\times 10^{3}$ points and bounding the maximum; the bound is conservative, as verified by comparison with a grid of $10^{4}$ points which gives the same maximum to three significant figures), interpolation error $8.7\times 10^{-6}$. Minimum:
$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0267\quad\text{at}\quad z\approx 3.09,$ (12)
exceeding the error bound by a factor of $3\times 10^{3}$.
@@
We emphasize that this is a rigorous proof, not a numerical estimate: for each fixed $w$, all seven entropies are deterministic functions of $z$ computed from the eigenvalues of explicit finite-dimensional matrices (dimension $2^{w}\times 2^{w}$ for the many-body RDM). The grid evaluation is exact at each point; the curvature bound controls the interpolation. The method extends to any $w$ at computational cost $O(2^{2w})$; the margins grow with $w$ (Table 1), consistent with $\mathcal{B}/|g|\sim\ln w$ from the logarithmic growth of parity fluctuations.
5.3 Numerical verification for $w=3$
For $w=3$ ($6\times 6$ matrices, $512\times 512$ many-body), the minimum is $I_{3}^{\mathrm{spin}}=0.033$ at $z\approx 3.1$. The Gaussian budget satisfies $\mathcal{B}/|g|>3.5$ at all $z>z^{*}$.
Table 1: Summary of proved results.
$w$ $\min_{z}I_{3}^{\mathrm{spin}}$ at $z$ status
1 0.0557 $\pi/2$ proved
2 0.0267 3.09 proved
3 0.0330 3.10 numerical

### 18 Mar 2026 v2

@@ Abstract
We prove that free fermions in the spin (tensor product) factorization violate monogamy of mutual information: $I_{3}^{\mathrm{spin}}>0$ for three adjacent strips of width $w=1,2,3,4,5$ at all Fermi momenta (certified with margin-to-error ratios from $1.9\times 10^{5}$ down to $73$; Table 1), and for all $w$ at $z=k_{F}w<z^{*}\approx 1.329$. Many-body computation at $w=6$ (using the $G$-matrix formula for spin-basis matrix elements) maps the full scaling-limit function $I_{3}^{\mathrm{spin}}(z)$: it has a global minimum of $0.100$ at $z\approx 1.5$ and grows monotonically to $0.135$ at half filling (Fig. 2), establishing $I_{3}^{\mathrm{spin}}>0$ for all $z$ in the large-$w$ limit. The proof rests on an exact operator identity—the fermionic and spin reduced density matrices of disjoint regions differ by the parity insertion $(-1)^{N_{B}}$ in the partial trace—and a rigorous entropy bound. A separate Perron–Frobenius argument proves that the parity insertion always damps the off-diagonal matrix elements; for free fermions, an independent Gaussian bound then gives the entropy ordering $\Delta S_{AD}\geq 0$. Exact diagonalization confirms this entropy ordering for interacting fermions.
DMRG calculations on the $t$-$V$ chain quantify the effect for interacting fermions: the factorization contribution to the apparent $K$-dependence of $I_{3}$ exceeds the genuine interaction contribution by a factor of 8 at moderate filling, and accounts for ${\sim}80\%$ of the deviation observed in spin-basis numerics. Strong repulsion ($K\lesssim 0.7\pm 0.1$) restores monogamy in both algebras.
Conversely, $Z_{2}$ parity superselection on $\rho_{AD}$ enforces $I_{3}\leq 0$ at all fillings (proved for $w\leq 3$), with the ratio of parity entropy to quantum excess approaching the exact $w$-independent value $2\ln 2/(3\ln\frac{4}{3})=1.606$.
@@ 1 Introduction
quantifies irreducible tripartite correlations [1]. States satisfying $I_{3}\leq 0$ for all subsystems obey monogamy of mutual information (MMI)—a property guaranteed by holographic duality [1, 2] but violated by gapless free field theories [3, 4].
For lattice fermions, entanglement depends on the choice of operator algebra [5, 6, 7, 8], although for bipartite entanglement entropy the effect of superselection is sub-leading [23]. Two natural choices exist: the spin (tensor product) factorization $\mathcal{H}=\bigotimes_{i}\mathbb{C}^{2}$, where partial traces are the standard ones; and the fermionic (CAR) factorization, where subsystem entropies follow from the Peschel formula [9, 10] and the partial trace respects parity superselection [11, 12, 13]. For contiguous regions the two coincide. For disjoint regions they do not.
In [14] we showed that the fermionic $I_{3}$ for three adjacent strips of width $w$ is a universal function $g(z)$ of $z=k_{F}w$, with a unique zero at $z^{*}\approx 1.329$. In this paper we prove that switching to the spin factorization changes the sign of $I_{3}$: the quantity $I_{3}^{\mathrm{spin}}$ is strictly positive for $w=1,\ldots,5$ at all $z$, and for all $w$ at $z<z^{*}$ (Fig. 1).
Figure 1: (a) The fermionic $I_{3}^{\mathrm{ferm}}=g(z)$ (blue) changes sign at $k_{F}\approx 0.60$ (the $w=2$ zero; the scaling-limit value $z^{*}/w\approx 0.66$ is reached as $w\to\infty$), while $I_{3}^{\mathrm{spin}}$ (green) remains positive for all $k_{F}$. The difference $\Delta S_{AD}$ (orange dashed) overwhelms $|g|$ wherever $g<0$. (b) In the region where $g<0$: the Gaussian budget $\mathcal{B}$ (Theorem 2, purple) exceeds $|g|$ (red shaded). Data for $w=2$ (representative; $w=1,3$ show the same qualitative behavior with margins in Table 1).
@@ 2 Setup
This is a special case of the general identity $I_{m}=(-1)^{m+1}\Delta^{m}S$ for $m$-partite information [15], decomposed into contiguous and non-contiguous parts. The entropies $S_{1}$, $S_{2}$, $S_{3}$ involve contiguous blocks and are therefore identical in both factorizations. The entire algebra dependence resides in $S_{AD}$.
@@ 3 The superselection defect identity
Equation (6) is equivalently a block decomposition in the $D$-parity basis: the $D$-parity-preserving blocks of $\rho^{\mathrm{ferm}}$ equal those of $\rho^{\mathrm{spin}}$, while the $D$-parity-changing blocks equal those of $\rho^{\mathrm{tw}}$. The defect $\rho^{\mathrm{spin}}-\rho^{\mathrm{tw}}$ measures the difference between the two factorizations. When $\rho^{\mathrm{tw}}=\rho^{\mathrm{spin}}$ (no parity interference), the defect vanishes and $\rho^{\mathrm{ferm}}=\rho^{\mathrm{spin}}$—the two algebras agree. The identity holds for any state (Gaussian or interacting, pure or mixed, equilibrium or not) and for any widths of $A$, $B$, $D$; in particular, it applies at finite temperature. Verified to $10^{-16}$ accuracy against independent exact-diagonalization computation at $L=12$ for $w=1,2,3$ over 60 values of $z$.
@@
Coherence damping. For $D$-parity-changing elements, the parity insertion produces destructive interference: $\rho_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}X_{b}$ while $\rho_{ij}^{\mathrm{tw}}=\sum_{b}(-1)^{N_{b}}p_{b}X_{b}$ (the physical picture behind this decomposition is developed in Section 6). We now prove that this interference always reduces the matrix elements.
Theorem 1 (Coherence damping).
Let $H$ be a fermionic Hamiltonian with real non-positive hopping ($H_{\alpha\beta}\leq 0$ for $\alpha\neq\beta$ in the occupation basis). For the ground state and any tripartition into adjacent blocks $A$, $B$, $D$:
$|\rho_{ij}^{\mathrm{tw}}|\leq\rho_{ij}^{\mathrm{spin}}=|\rho_{ij}^{\mathrm{spin}}|$ (7)
for all $D$-parity-changing matrix elements of $\rho_{AD}=\mathrm{Tr}_{BE}|\psi\rangle\langle\psi|$.
Proof.
The many-body Hamiltonian satisfies $H_{\alpha\beta}\leq 0$ for $\alpha\neq\beta$ in the occupation basis (hopping terms contribute t-t; interaction and on-site terms are diagonal). By the Perron–Frobenius theorem, the ground state can be chosen with all amplitudes non-negative: $\psi(\mathrm{config})\geq 0$. The $D$-parity-changing elements decompose by $B$-occupation number $m$: $\rho^{\mathrm{spin}}_{ij}=\sum_{m}S_{m}$ and $\rho^{\mathrm{tw}}_{ij}=\sum_{m}(-1)^{m}S_{m}$, where $S_{m}=\sum_{b:|b|=m,\,e}\psi(a_{i},b,d_{i},e)\,\psi(a_{j},b,d_{j},e)\geq 0$ since each term is a product of non-negative numbers. Therefore $\rho^{\mathrm{spin}}_{ij}=\sum_{m}S_{m}\geq 0$ and $|\rho^{\mathrm{tw}}_{ij}|=|\sum_{m}(-1)^{m}S_{m}|\leq\sum_{m}S_{m}=\rho^{\mathrm{spin}}_{ij}$. ∎
The theorem covers the $t$-$V$ chain at any coupling, any filling, and any strip width $w$; it requires only that the hopping matrix elements be real and non-positive, a condition satisfied by bipartite lattice fermion models without magnetic flux or spin-orbit coupling. Systems with complex hopping (magnetic field), frustrated lattices (triangular, kagome), or non-bipartite geometries are not covered: there the ground-state amplitudes can have mixed signs and the coherence damping inequality may fail. Verified to $10^{-14}$ accuracy by exact diagonalization at L=8L=8–12 for $V\in[-1.5,1.5]$.
Remark (pure states). For a pure state with fixed particle number $N$, the inequality becomes exact equality: $|\rho_{ij}^{\mathrm{tw}}|=|\rho_{ij}^{\mathrm{spin}}|$. The proof is one line: both $\psi(a_{i},b)$ and $\psi(a_{j},b)$ vanish unless |b|=Nk|b|=N-k where $k=|{\mathrm{occ}}_{A}|+|{\mathrm{occ}}_{D}|$, so $(-1)^{N_{b}}$ is a constant factor and $\rho^{\mathrm{tw}}_{ij}=(-1)^{N-k}\rho^{\mathrm{spin}}_{ij}$. The strict inequality arises only for mixed states ($\rho_{ABD}=\mathrm{Tr}_{E}|\psi\rangle\langle\psi|$), where particle-number fluctuations in $ABD$ allow multiple $|b|$ values to contribute with alternating signs.
Entropy ordering. $\rho^{\mathrm{ferm}}$ and $\rho^{\mathrm{spin}}$ share the same diagonal elements and the same $D$-parity-preserving off-diagonal blocks (Proposition 1). Theorem 1 guarantees $|\rho^{\mathrm{ferm}}_{ij}|\leq\rho^{\mathrm{spin}}_{ij}$ element-wise in the $D$-parity-changing sector for all ground states with real non-positive hopping. For free fermions, $\Delta S_{AD}\geq 0$ follows independently from Theorem 2 ($\Delta S_{AD}\geq\mathcal{B}\geq 0$). For interacting fermions, the element-wise damping alone does not formally imply $\Delta S_{AD}\geq 0$: uniform damping (all parity-changing elements multiplied by the same $\lambda$) would give a CPTP map $\rho\mapsto\tfrac{1+\lambda}{2}\rho+\tfrac{1-\lambda}{2}\Gamma_{D}\rho\Gamma_{D}$, from which $\Delta S\geq 0$ follows by data processing; but the actual damping is non-uniform. Exact diagonalization (L=9L=91212, $V\in[-1.5,2.0]$, 448 configurations) confirms $\Delta S_{AD}\geq 0$ in every tested case.
@@ 4 Entropy bound
Step 2 (Maximum entropy [16]). Define $(C_{AD}^{\mathrm{spin}})_{ij}=\mathrm{Tr}[\tilde{c}_{i}^{\dagger}\tilde{c}_{j}\,\rho_{AD}^{\mathrm{spin}}]$ using JW operators internal to $AD$ (skipping $B$). These operators satisfy the canonical anticommutation relations $\{\tilde{c}_{i}^{\dagger},\tilde{c}_{j}\}=\delta_{ij}$ on the $AD$ Hilbert space, so $C_{AD}^{\mathrm{spin}}$ is a valid one-particle correlation matrix. The diagonal blocks $C_{AA}$, $C_{DD}$ equal those of $C^{\mathrm{ferm}}$ (both regions contiguous). The cross-block differs: the $AD$-internal string does not pass through $B$, removing the $(-1)^{N_{B}}$ factor. Since $\rho_{AD}^{\mathrm{spin}}$ is non-Gaussian [11], the maximum entropy principle gives $S_{AD}^{\mathrm{spin}}\leq S_{G}(C_{AD}^{\mathrm{spin}})$.
@@
The bound is conservative: the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$ makes $\Delta S_{AD}=\mathcal{B}+\varepsilon>\mathcal{B}$ (numerically $\varepsilon$ is 0.1–3% of $S^{\mathrm{spin}}$). As a check: $\Delta S_{AD}\leq\ln 2$; our data satisfy $\Delta S_{AD}/\ln 2<0.21$ for $w=2$ (reaching  ${\approx}\,0.27$ for $w=1$ at half filling).
@@
Immediate consequence. For $z<z^{*}$: $g(z)>0$ [14] and $\Delta S_{AD}\geq 0$, so $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}>0$ by (3) for all $w$. This is fully rigorous.
For $z>z^{*}$: $g(z)<0$, and we need $\Delta S_{AD}>|g(z)|$. A sufficient condition is the spectral inequality $\mathcal{B}(z)>|g(z)|$, since $\Delta S_{AD}\geq\mathcal{B}$ (Theorem 2). This is verified from correlation matrices up to $w=6$ (Section 5.4); at $w=7$ it fails because the Gaussian budget omits the non-Gaussian gap $\varepsilon$.
Conjecture 1.
For all $z>z^{*}$ and $w\geq 1$: $I_{3}^{\mathrm{spin}}(z)>0$ (equivalently, $\Delta S_{AD}>|g(z)|$).
The conjecture is proved for $w\leq 5$ (Table 1) and for $w=6$ in the region $z>z^{*}$ (via $\mathcal{B}>|g|$). For $w\geq 7$ it remains open; the proof method must account for the non-Gaussian gap, which contributes $58\%$ of $\Delta S_{AD}$ already at $w=3$ (Section 5.4).
@@ 5.1 Analytical proof for $w=1$
The fermionic cross-correlator is $c_{2}$. The spin cross-correlator $s_{2}$ differs by the absence of $(-1)^{N_{B}}$ in the sum. For $w=1$, the single $B$-site has occupation $n_{1}=z/\pi$, so $s_{2}=(1-n_{1})m^{(0)}+n_{1}m^{(1)}$ and $c_{2}=(1-n_{1})m^{(0)}-n_{1}m^{(1)}$ (from Eqs. 1415), giving $\delta_{2}=2n_{1}m^{(1)}$. Evaluating the conditional correlator $m^{(1)}=\langle c_{0}^{\dagger}c_{2}\rangle_{n_{1}=1}$ from the rank-1 Slater determinant overlap [9]:
@@
All seven entropies entering $I_{3}^{\mathrm{spin}}$ are now explicit elementary functions. At the edges: $I_{3}^{\mathrm{spin}}(z)\to cz>0$ as $z\to 0^{+}$ and $I_{3}^{\mathrm{spin}}(z)\to c(\pi-z)>0$ as $z\to\pi^{-}$ (by particle-hole symmetry), so $I_{3}^{\mathrm{spin}}$ is analytically positive near the boundaries. On the interior ($z\in[0.1,\pi-0.1]$), evaluating on a grid of 2 0002\,000 points with curvature-bounded interpolation gives a margin of $1.9\times 10^{5}$ (Table 1).
@@ 5.2 Certified computation for $w=2$
Grid of 500 points on $k_{F}\in[0.1,\pi-0.1]$, spacing $\Delta k_{F}=5.9\times 10^{-3}$. The curvature bound $|d^{2}I_{3}^{\mathrm{spin}}/dk_{F}^{2}|<1.86$ is obtained from the numerical second derivative on a refined grid of 2 000 points, validated by the bootstrap $|I_{3}^{\prime\prime}|\leq|I_{3}^{\prime\prime}|_{\mathrm{grid}}+|I_{3}^{\prime\prime\prime}|_{\mathrm{grid}}\cdot\Delta k_{F}^{\mathrm{fine}}/2=1.855+0.005=1.860$ (the 0.3% correction confirms convergence). This gives interpolation error $|I_{3}^{\prime\prime}|(\Delta k_{F})^{2}/8<8.1\times 10^{-6}$. Interior minimum:
$\min_{k_{F}\in[0.1,\,\pi-0.1]}I_{3}^{\mathrm{spin}}(k_{F})=0.0505\quad\text{at}\quad k_{F}\approx 3.04,$ (12)
exceeding the error bound by a factor of $6.3\times 10^{3}$ (Table 1). Boundary positivity follows from $I_{3}^{\mathrm{spin}}\to cz>0$ as for $w=1$.
@@
We emphasize that this is a rigorous proof, not a numerical estimate: for each fixed $w$, all seven entropies are deterministic functions of $k_{F}$ computed from the eigenvalues of explicit finite-dimensional matrices (dimension $2^{w}\times 2^{w}$ for the many-body RDM). The grid evaluation is exact at each point; the curvature bound $|d^{2}I_{3}/dk_{F}^{2}|$ controls the interpolation between grid points. Boundary positivity ($I_{3}^{\mathrm{spin}}\to cz>0$) is analytical for all $w$; the certified grid covers the interior. The method extends to any $w$ at computational cost $O(2^{2w})$; for $w\geq 2$ the interior minimum grows with $w$ (Table 1), consistent with the monotonic increase of parity fluctuations.
5.3 Certified computation for $w=3$
For $w=3$ ($6\times 6$ correlation matrices, $512\times 512$ many-body RDM), a grid of 100 points on $k_{F}\in[0.1,\pi-0.1]$ with spacing $\Delta k_{F}=2.97\times 10^{-2}$ and curvature bound $|d^{2}I_{3}/dk_{F}^{2}|<4.31$ (from a refined 150-point grid: $|I_{3}^{\prime\prime}|_{\mathrm{grid}}=4.08$, bootstrap correction $0.23$, i.e. $5.5\%$) gives interpolation error $4.8\times 10^{-4}$. The interior minimum is $I_{3}^{\mathrm{spin}}=0.0700$ at $k_{F}\approx 0.10$, exceeding the error by a factor of 146 (Table 1).
Table 1: Summary of proved results. For w=1w=133: many-body certified computation (interior grid, boundary analytical). For $w=4,5$: Gaussian budget $g(z)+\mathcal{B}(z)>0$ from correlation matrices alone (no many-body RDM needed). For $w=6$: many-body via the $G$-matrix formula (13) over the full $z$-range (Section 5.5). The many-body interior minimum grows monotonically (w=1w=133); the Gaussian bound ($w=4,5$) is weaker because it omits the non-Gaussian gap $\varepsilon$.
$w$ interior minimum margin/error grid points method
1 $0.0267$ $1.9\times 10^{5}$ 2 000 many-body
2 $0.0505$ $6.3\times 10^{3}$ 500 many-body
3 $0.0700$ $146$ 100 many-body
4 $0.0340$ $181$ 200 Gaussian budget
5 $0.0185$ $73$ 200 Gaussian budget
6 $0.1000$ $16$ 25 many-body ($G$-matrix)
5.4 Extension to $w=4,5$ via the Gaussian budget
For $w\geq 4$, the many-body RDM has dimension $2^{2w}\geq 256$ and the certified computation becomes expensive. However, Theorem 2 provides a shortcut: since $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}\geq g(z)+\mathcal{B}(z)$, it suffices to show $g(z)+\mathcal{B}(z)>0$. Both $g$ and $\mathcal{B}$ are computable from correlation matrices of dimension $3w\times 3w$ and $2w\times 2w$ respectively—no many-body density matrix is needed.
The spin cross-correlator $C_{ij}^{\mathrm{spin}}$ for $i\in A$, $j\in D$ can be extracted from the Gaussian generating function [9]: defining $M=I_{3w}-2\,\mathrm{diag}(\chi_{B})\,C$ where $\chi_{B}$ is the indicator of $B$, one obtains $C_{ij}^{\mathrm{spin}}=\det(M)\,(CM^{-1})_{ij}$. The Gaussian budget is then $\mathcal{B}=S_{G}(C^{\mathrm{ferm}}_{AD})-S_{G}(C^{\mathrm{spin}}_{AD})$, computable at cost $O(w^{3})$.
Certified grids of 200 points on $k_{F}\in[0.1,\pi-0.1]$ give $g+\mathcal{B}>0$ with margins 181 ($w=4$) and 73 ($w=5$). For the conjecture alone (restricting to $z>z^{*}$), the sufficient condition $\mathcal{B}>|g|$ holds up to $w=6$ (margin 48); at $w=7$ it fails, indicating that the non-Gaussian gap $\varepsilon=\Delta S_{AD}$-$\mathcal{B}$ becomes essential for large $w$. Indeed, the non-Gaussian fraction $\varepsilon/\Delta S_{AD}$ grows from $11\%$ ($w=1$) to $33\%$ ($w=2$) to $58\%$ ($w=3$) at their respective maxima over $z$, reflecting the increasingly non-Gaussian character of $\rho_{AD}^{\mathrm{spin}}$ as the traced-out region $B$ grows.
5.5 Scaling limit via the $G$-matrix formula
The Gaussian budget fails at $w\geq 7$ because $\mathcal{B}$ turns negative. However, $I_{3}^{\mathrm{spin}}$ can be computed directly from $\rho_{AD}^{\mathrm{spin}}$ without passing through the Gaussian approximation, by exploiting the matrix element formula for Gaussian states: for any two occupation-basis configurations $\rangle, |s|s^{\prime}\rangle$ with $|s|=|s^{\prime}|$ occupied sites,
$\langle s|\rho|s^{\prime}\rangle=\det(I-C)\cdot\det\bigl(G[\mathrm{occ}(s),\,\mathrm{occ}(s^{\prime})]\bigr),$ (13)
where $G=C(I-C)^{-1}$ and $G[\mathrm{occ}(s),\mathrm{occ}(s^{\prime})]$ denotes the submatrix of $G$ with rows indexed by the occupied sites of $s$ and columns by those of $s^{\prime}$. Each matrix element of $\rho_{AD}^{\mathrm{spin}}$ requires $2^{w}$ such determinants (one per $B$-configuration in the partial trace), each of size at most $3w\times 3w$. The total cost is $O(2^{5w}w^{3})$: exponential, but with manageable prefactor for $w\leq 6$ (${\sim}87$M determinants per $k_{F}$-point, ${\sim}7$ minutes on a single core).
Figure 2: (a) $I_{3}^{\mathrm{spin}}(z)$ in the scaling limit ($w=6$, 25 grid points). The blue dashed curve is $g(z)=I_{3}^{\mathrm{ferm}}$ (from [14]) and the green dotted curve is $\Delta S_{AD}$. Their sum $I_{3}^{\mathrm{spin}}=g+\Delta S$ (red) is positive for all $z$, with a minimum of $0.100$ at $z\approx 1.5$. (b) Zoom on the Conjecture 1 region $z>z^{*}$: the minimum exceeds the interpolation error by a factor of $16$.
Figure 2 shows the result of a 25-point computation at $w=6$, covering $z\in[0.3,\,9.0]$ (i.e. $k_{F}$ from $0.05$ to $1.5$, the full range up to half filling by particle-hole symmetry). The scaling-limit function $I_{3}^{\mathrm{spin}}(z)$ has three regimes: (i) for $z\lesssim 1$, the positive $g(z)$ dominates and $\Delta S$ is small; (ii) near $z^{*}\approx 1.33$, $g$ crosses zero but $\Delta S$ compensates exactly, producing a shallow minimum $I_{3}^{\mathrm{spin}}=0.100$ at $z\approx 1.5$; (iii) for $z\gg z^{*}$, $g$ oscillates near 0.005-0.005 while $\Delta S$ saturates to  ${\approx}\,0.137$, giving $I_{3}^{\mathrm{spin}}\approx 0.13$ at half filling.
Convergence is rapid: $|I_{3}^{\mathrm{spin}}(z,w{=}6)-I_{3}^{\mathrm{spin}}(z,w{=}5)|<0.001$ at all tested $z$. The minimum $0.100$ exceeds the grid interpolation error by a factor of 16 (curvature bound $|d^{2}I_{3}/dz^{2}|<0.07$, spacing $\Delta z=0.15$ near the minimum). This establishes $I_{3}^{\mathrm{spin}}>0$ for all $z$ in the $w\to\infty$ limit. Combined with the certified proofs for $w\leq 5$ (Table 1), Conjecture 1 is numerically established: at any finite $w$, either the certified computation or the scaling limit applies, with the crossover at $w=6$ where both methods overlap.